The length of a rectangular garden is 1 ft less than 2 times its width. If the perimeter of the garden is 46 ft, find the dimensions of the garden.
Answer by pwac(253)

Let width = x
so length=2x-1
Perimeter is x+x+2x-1+2x-1=46 Do sum on left:
6x-2=46 Add 2 to both sides
6x=48 Divide both sides by 6:
x=8
So width=8ft.
Length=2times8-1=15ft.
Pete